# Simplifying FFT-based methods for mechanics with automatic differentiation

### v1

@@ Abstract
Fast-Fourier Transform (FFT) methods have been widely used in solid mechanics to address complex homogenization problems. However, current FFT-based methods face challenges that limit their applicability to intricate material models or complex mechanical problems. These challenges include the manual implementation of constitutive laws and the use of computationally expensive and complex algorithms to couple microscale mechanisms to macroscale material behavior. Here, we incorporate automatic differentiation (AD) within the FFT framework to mitigate these challenges. We demonstrate that AD-enhanced FFT-based methods can derive stress and tangent stiffness directly from energy density functionals, facilitating the extension of FFT-based methods to more intricate material models. Additionally, AD simplifies the calculation of homogenized tangent stiffness for microstructures with complex architectures and constitutive properties. This enhancement renders current FFT-based methods more modular, enabling them to tackle homogenization in complex multiscale systems, especially those involving multiphysics processes. Our work will simplify the numerical implementation of FFT-based methods for complex solid mechanics problems.
@@ 1 Introduction
A crucial aspect of computational mechanics is understanding how microscale mechanisms impact macroscale material behavior. Features such as the distribution of heterogeneities, the formation of structures at the micro level, and physical processes occurring on the microscale can significantly influence the properties homogenized at a larger scale. The Fast-Fourier Transform (FFT) methods [1, 2, 3] are particularly well-suited for tackling such ng homogenization problems due to their computational efficiency [4, 5], ease of implementation [6], and low memory footprint. Nonetheless, certain aspects of FFT frameworks limit their quick adaptation to various applications and extension to intricate mechanical systems.
@@
A significant challenge with FFT-based methods is that they still rely on manually implemented stress-strain relationships, which are computed “by hand” either analytically or through numerical differentiation [4]. This makes the implementation of complex constitutive relations (e.g., hyperelastic materials and elastoplastic materials) intricate and error-prone within the FFT framework, especially for cases involving multi-physical processes [7]. Hence, one of their major strength (i.e. ease of implementation) is lost for more complex problems. Another prominent limitation lies in determining the homogenized tangent stiffness in multiscale simulations [8, 9, 10, 11]. Current FFT-based multiscale frameworks require auxiliary solution schemes [12] to couple microscale mechanisms to the macroscale through a constitutive tangent stiffness, which can be computationally expensive and inaccurate for complex material models and large computational domains.
@@
Other numerical methods have similar challenges, which have been overcome by adopting a differentiable framework, which automates the process of computing derivatives [13]. This allows the formulation of the governing equation in a high-level manner (usually as an expression of the functional energy), and the complex Jacobian and the Hessian are then strictly derived through automatic differentiation (AD) up to machine precision [7, 14]. Among prominent examples of such differentiable frameworks within the field of solid mechanics is FEniCSx [15] for Finite Element Methods. Concurrently, with the rise of automatic differentiation (AD) libraries such as pyTorch [16] and JAX [17], the differentiable framework has become quite appealing, which is evident by the shift in porting existing frameworks such as molecular dynamics, discrete element method, and finite element method [18, 19, 20, 21, 22] to a differentiation-based framework. Similarly, AD presents great potential for FFT-based methods [23] but has yet to be developed in a general and ease-of-use manner to overcome the main challenges of FFT-based methods.
@@
In this (short) paper, we enhance the FFT framework by incorporating automatic differentiation. We investigate the capabilities of an AD-enhanced FFT framework and examine how automatic differentiation mitigates some of the aforementioned issues. Thereby making FFT-based frameworks more modular and easy to implement for complex mechanical problems.
@@ 2 Differentiable FFT-framework for solid mechanics
In this study, we adopt the Fourier-Galerkin spectral method [2, 3], and do so for several reasons. Primarily, the variational structure of this approach retains the convexity of the energy density [3, 4], which ensures that a unique solution to a given problem exists. Secondly, to guarantee that the solution produces a compatible strain state, the method employs a projection operator $\mathbb{G}$, which converts an arbitrary strain state to a compatible state (periodic and curl-free). This operator is independent of the material model and remains consistent for both small-strain and finite deformation scenarios [6]. Finally, unlike other methods [1], the Fourier-Galerkin method does not rely on a reference medium stiffness tensor, which can impact the convergence in nonlinear problems [6]. We will show that these characteristics of the Fourier-Galerkin approach make integrating automatic differentiation within the FFT-based framework straightforward.
@@
Here, we will show how automatic differentiation can be applied in an FFT-based framework. We will describe three use cases of AD, namely the stress computation in elastic and hyperelastic materials, the computation of consistent tangent operator for nonlinear problems, and the computation of homogenized stiffness operator for multiscale simulations. For all cases, we include both the arithmetic operations as well as the counterpart pseudo-code to highlight the general applicability and ease of use of AD in FFT-based methods. To this end, we employ the open-source JAX library [17] for automatic differentiation and demonstrate that only a few lines of code are needed to implement a general solution.
First, we consider the stress computation of elastic and hyperelastic materials, for which the stress function $\bm{\sigma}$ is the first derivative of the strain energy density, as given by
@@
The application of automatic differentiation on Eq. 2 using the JAX library [17] is straightforward: {minted}[frame=lines, framesep=1mm, baselinestretch=0.7, fontsize=, ]python def strain_energy(strain): # compute the energy based on the input ‘strain‘ return energy
@@
# create a function to compute stresses using the # derivative of energy w.r.t ‘strain‘ (Eq 2) compute_stresses = jax.jacfwd(strain_energy)
@@
# using the above created function to compute stresses sigma = compute_stresses(strain) Here, we define a function that receives microscale strain as input and outputs the corresponding energy density. A function to calculate stresses can then be derived using the forward differentiation technique [13, 17]. We note that the input strain to the strain-energy function may be either a small strain ($\bm{\varepsilon}$) or a deformation gradient ($\bm{F}$) based on the defined energy density. In Sec. 3.1, we will present examples for various material models under small-strain and finite deformation settings.
@@
Next, we integrate automatic differentiation to ease the numerical implementation of the nonlinear form of Equation 1. In standard practice, Equation 1 is linearized, and the Newton-Raphson method is used to find the equilibrium. The linearization process involves computing the incremental stresses and the consistent tangent operator at each step of the Newton-Raphson scheme, which is expressed as follows,
@@
$\delta\bm{\sigma}=-\underbrace{\dfrac{\partial\bm{\sigma}(\bm{e})}{\partial\bm{e}}}_{\mathbb{K}}\delta\bm{e}\leavevmode\nobreak\ .$ (3)
Consequently, the local tangent stiffness $\mathbb{K}$ is needed at the previous strain for the computation of the incremental stresses. With automatic differentiation, the incremental stress can be computed as follows {minted}[frame=lines, framesep=1mm, baselinestretch=0.7, fontsize=, ]python # create a function to compute stresses compute_stresses = jax.jacfwd(strain_energy)
# computing the incremental stresses as given in Eq 3 dsigma = jax.jvp(compute_stresses, (prev_strain,), (dstrain,))[1] Here, we utilize a JAX functionality to represent the local tangent stiffness as a push-forward function (or a linear operator) that maps incremental strains $\delta\bm{e}$ lying in the tangent space of $\bm{e}$ to the incremental stresses $\delta\bm{\sigma}$ lying in the tangent space of the stress domain. This facilitates the direct computation of incremental stress as a Jacobian-vector product [25, 26]. Furthermore, representing $\mathbb{K}$ as a push-forward map instead of explicitly storing it in a matrix form is expected to reduce the memory footprint of FFT-based methods significantly.

### v2

# Simplifying FFT-based methods for solid mechanics with automatic differentiation

@@ Abstract
Fast-Fourier Transform (FFT) methods have been widely used in solid mechanics to address complex homogenization problems. However, current FFT-based methods face challenges that limit their applicability to intricate material models or complex mechanical problems. These challenges include the manual implementation of constitutive laws and the use of computationally expensive and complex algorithms to couple microscale mechanisms to macroscale material behavior. Here, we incorporate automatic differentiation (AD) within the FFT framework to mitigate these challenges. We demonstrate that AD-enhanced FFT-based methods can derive stress and tangent stiffness directly from energy density functionals, facilitating the extension of FFT-based methods to more intricate material models. Additionally, automatic differentiation simplifies the calculation of homogenized tangent stiffness for microstructures with complex architectures and constitutive properties. This enhancement renders current FFT-based methods more modular, enabling them to tackle homogenization in complex multiscale systems, especially those involving multiphysics processes. Furthermore, we illustrate the use of the AD-enhanced FFT method for problems that extend beyond homogenization, such as uncertainty quantification and topology optimization where automatic differentiation simplifies the computation of sensitivities. Our work will simplify the numerical implementation of FFT-based methods for complex solid mechanics problems.
@@ 1 Introduction
A crucial aspect of computational multiscale mechanics is understanding how microscale mechanisms impact macroscale material behavior. Features such as the distribution of heterogeneities, the formation of structures at the micro level, and physical processes occurring on the microscale can significantly influence the properties homogenized at a larger scale. The Fast-Fourier Transform (FFT) methods [1, 2, 3, 4] are particularly well-suited for tackling such ng homogenization problems due to their computational efficiency [5, 6], ease of implementation [7], and low memory footprint. Nonetheless, certain aspects of FFT frameworks limit their quick adaptation to various applications and extension to intricate mechanical systems.
@@
A significant challenge with FFT-based methods is that they still rely on manually implemented stress-strain relationships, which are computed “by hand” either analytically or through numerical differentiation [5]. This makes the implementation of complex constitutive relations (e.g., hyperelastic materials and elastoplastic materials) intricate and error-prone within the FFT framework, especially for cases involving multi-physical processes [8]. Hence, one of their major strength (i.e. ease of implementation) is lost for more complex problems. Another prominent limitation lies in determining the homogenized tangent stiffness in multiscale simulations [9, 10, 11, 12, 13]. Current FFT-based multiscale frameworks require auxiliary solution schemes [14] to couple microscale mechanisms to the macroscale through a constitutive tangent stiffness, which can be computationally expensive and inaccurate for complex material models and large computational domains.
@@
Other numerical methods have similar challenges, which have been overcome by adopting a differentiable framework, which automates the process of computing derivatives [15]. This allows the formulation of the governing equation in a high-level manner (usually as an expression of the functional energy), and the complex Jacobian and the Hessian are then strictly derived through automatic differentiation (AD) up to machine precision [8, 16]. Among prominent examples of such differentiable frameworks within the field of solid mechanics is FEniCSx [17] for Finite Element Methods. Concurrently, with the rise of AD libraries such as PyTorch [18] and JAX [19], the differentiable framework has become quite appealing, which is evident by the shift in porting existing frameworks such as molecular dynamics, discrete element method, and finite element method [20, 21, 22, 23, 24] to a differentiation-based framework. Similarly, automatic differentiation presents great potential for FFT-based methods [25] but has yet to be developed in a general and ease-of-use manner to overcome the main challenges of FFT-based methods.
@@
In this paper, we enhance the FFT framework by incorporating automatic differentiation. We investigate the capabilities of an AD-enhanced FFT framework and examine how automatic differentiation mitigates some of the aforementioned issues. Thereby making FFT-based frameworks more modular and easy to implement for complex solid mechanics problems.
2 How automatic differentiation works?
Consider a function $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ that maps an $n$-dimensional input $x$ to an $m$-dimensional output. To determine the derivatives of the outputs relative to the inputs, specifically $J_{ij}=\partial f_{i}/\partial x_{j}$ where $i\in m$ and $j\in n$, one typically depends on the hand-calculated closed-form expression of the derivative of the function. In case the function $f$ has a complex closed-form mathematical expression, one can use symbolic differentiation to get the derivatives [26]. In computational solid mechanics, $f$ may represent a complex function consisting of numerous procedures that do not possess an explicit mathematical expression. For example, $f$ might be a function that computes the Fourier transform of the input data with a conditional statement (if-then-else) or operates a minimization solver with for-loops on the inputs. In such cases, neither analytical nor symbolic differentiation is feasible, requiring the use of alternative approaches to estimate the derivatives $\partial f_{i}/\partial x_{j}$. A classical approach is numerical differentiation. Its results are very sensitive to the perturbations used in the calculations, which may result in inaccuracies or numerical stability issues. Alternatively, automatic differentiation overcomes such limitations by computing exact derivatives for such functions. This is achieved by decomposing a complex function $f$ into elementary functions, for which exact derivatives are known, and interconnecting them through basic arithmetic operations such as addition and subtraction. The chain rule of calculus is then applied to compute exact partial derivatives. For instance, if the function $f$ can be decomposed as $f=h(g(x))$, such that $x\in\mathbb{R}^{n},g:\mathbb{R}^{n}\to\mathbb{R}^{k}$ and $h:\mathbb{R}^{k}\to\mathbb{R}^{m}$, then the partial derivative $df_{i}/dx_{j}$ is computed as
$J_{ij}=\dfrac{\partial f_{i}}{\partial x_{j}}=\sum_{p=1}^{k}\dfrac{\partial h_{i}}{\partial g_{p}}\dfrac{\partial g_{p}}{\partial x_{j}}\leavevmode\nobreak\ .$ (1)
As automatic differentiation utilizes the exact derivatives of the sub-functions to determine the derivative of $f$, it is not affected by issues related to numerical stability or round-off errors. The derived values retain the same level of precision as the function results, thereby enabling the accurate calculation of derivatives even in the case of nonlinear or non-monotonic functions. Moreover, automatic differentiation can be repeatedly applied to a function to derive exact higher-order derivatives without accumulating errors.
Figure 1: Forward vs. Reverse mode for a composable function $f:g\circ h$ (a) Schematic illustration showing automatic differentiation applied to function $f:\mathbb{R}\to\mathbb{R}$. The left column shows the execution graph for computing output of $f$. The center column illustrates the execution graph for calculating the derivative of $f$ via forward-mode. Observe that each node representing a function call is substituted by its derivative and is traversed in the original function call direction. The right column indicates the call direction for the reverse-mode, note the backwards arrows. (b) Schematic illustration showing automatic differentiation applied to function $f:\mathbb{R}^{1}\to\mathbb{R}^{2}$. Notice that the forward-mode differentiation requires only a single pass to compute the 2 partial derivatives whereas reverse-mode requires 2 passes to compute the partial derivatives. The 2 passes are indicated by the 2 separate graphs. (c) Schematic illustration showing automatic differentiation applied to function $f:\mathbb{R}^{2}\to\mathbb{R}^{1}$. Here the forward-mode requires 2 passes to compute the partial derivatives with respect 2 input values whereas reverse-mode requires a single pass.
Automatic differentiation has two ways to calculate the derivative of $f$ using the chain rule [15, 19]. The first technique involves evaluating the derivatives of sub-functions (inner derivatives) simultaneously with the function evaluation. This approach is known as forward-mode differentiation. Figure 1a illustrates forward-mode differentiation applied to a composite function $f=g\circ h$, mapping a scalar to a scalar, which requires only one pass to compute both the output of the function and its gradient. Alternatively, the second method involves first evaluating the function’s output and then moving backwards to determine the derivatives, starting with outer derivatives followed by inner ones. This is called reverse-mode differentiation and requires at least two evaluations to extract the function’s value and gradient. Figure 1a depicts this backward propagation of the function to first compute outer derivatives and then inner ones. The choice of method depends on how the function $f$ maps the input space to the output space. We demonstrate this with two simple cases for a function $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$: (i) when $n<m$, i.e., $f=(g_{1}(h(x)),g_{2}(h(x)))$ as shown in Figure 1b, and (ii) when $n>m$, i.e., $f=g(h(x,y))$ as shown in Figure 1c. In the case of $n=1<2=m$ (see Figure 1b), forward-mode requires a single pass to calculate both partial derivatives ($\partial f_{1}/\partial x,\partial f_{2}/\partial x$), while reverse-mode needs two backward computations to determine these derivatives. Conversely, when $n=2>1=m$, forward-mode requires two evaluations: one using $x$ (assuming $y=0$) and another using $y$ (assuming $x=0$), for computing partial derivatives ($\partial f/\partial x,\partial f/\partial y$). However, reverse-mode achieves this with a single pass. Therefore, as a basic rule, for a function $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$, forward-mode is computationally efficient when mapping from a lower-dimensional space to a higher-dimensional one, i.e., when $n\ll m$. In contrast, reverse-mode is efficient when the function maps from a higher to a lower dimension space, indicating $n\gg m$. We use this guideline throughout the paper to find the most efficient way to differentiate.
For this study, we use JAX for automatic differentiation due to its effective support for both forward- and reverse-mode differentiation, just-in-time (JIT) compilation, and efficient parallelization on GPUs. JAX is particularly well suited for scientific computing tasks requiring high-order derivatives, offering both flexibility and high performance.
@@ 3 Differentiable FFT-framework for solid mechanics
In this study, we adopt the Fourier-Galerkin spectral method [3, 4], and do so for several reasons. Primarily, the variational structure of this approach retains the convexity of the energy density [4, 5], which ensures that a unique solution to a given problem exists. Secondly, to guarantee that the solution produces a compatible strain state, the method employs a projection operator $\mathbb{G}$, which converts an arbitrary strain state to a compatible state (periodic and curl-free). This operator is independent of the material model and remains consistent for both small-strain and finite deformation scenarios [7]. We will show that these characteristics of the Fourier-Galerkin approach make integrating automatic differentiation within the FFT-based framework straightforward.
@@
Here, we will show how automatic differentiation can be applied in an FFT-based framework. We will describe three use cases of automatic differentiation, namely the stress computation in elastic and hyperelastic materials, the computation of consistent tangent operator for nonlinear problems, and the computation of homogenized stiffness operator for multiscale simulations. For all cases, we include both the arithmetic operations as well as the counterpart pseudo-code to highlight the general applicability and ease of use of automatic differentiation in FFT-based methods. To this end, we employ the open-source JAX library [19] for automatic differentiation and demonstrate that only a few lines of code are needed to implement a general solution.
3.1 Automatic computation of stress tensor from energy functional
We consider the stress computation of elastic and hyperelastic materials, for which the stress function $\bm{\sigma}$ is the first derivative of the strain energy density, as given by
@@ 3.1 Automatic computation of stress tensor from energy functional
The application of automatic differentiation on Eq. 3 using the JAX library [19] is straightforward as we show in Section 3.1. {listing} {minted}[frame=lines, framesep=1mm, baselinestretch=0.7, fontsize=, ]python def strain_energy(strain): # compute the energy based on the input ‘strain‘ return energy
@@
# create a function to compute stresses using the # derivative of energy w.r.t ‘strain‘ (Eq 2) compute_stresses = jax.jacrev(strain_energy)
@@
# using the above created function to compute stresses sigma = compute_stresses(strain) Automatic differentiation to compute stress from energy density functional Here, we define a function that receives microscale strain as input and outputs the corresponding energy density. Since the function maps a $2^{\mathrm{nd}}$-order strain tensor to a scalar value, we use the reverse-mode technique to differentiate the function to calculate stresses. We note that the input strain to the strain-energy function may be either a small strain ($\bm{\varepsilon}$) or a deformation gradient ($\bm{F}$) based on the defined energy density. In Sec. 4.1, we will present examples for various material models under small-strain and finite deformation settings.
3.2 Automatic computation of tangent moduli and incremental stresses
@@
We integrate automatic differentiation to ease the numerical implementation of the nonlinear Equation 2. In standard practice, Equation 2 is linearized, and the Newton-Raphson method is used to find the equilibrium. The linearization process involves computing the incremental stresses and the consistent tangent operator at each step of the Newton-Raphson scheme, which is expressed as follows,
@@
$\delta\bm{\sigma}=\underbrace{\dfrac{\partial\bm{\sigma}(\bm{e})}{\partial\bm{e}}}_{\mathbb{K}}\delta\bm{e}\leavevmode\nobreak\ .$ (4)
Consequently, the local tangent stiffness $\mathbb{K}$ is needed at the previous strain to calculate the incremental stresses. With automatic differentiation, there are three different ways in which one can differentiate the stress function to compute the local tangent stiffness. The straight-forward approach will be to simply differentiate the obtained stress function again. Since the stress function maps $2^{\mathrm{nd}}$-order strain tensor to $2^{\mathrm{nd}}$-order stress tensor, one can choose either the forward-mode or the reverse-mode for differentiation. Both will be computationally equivalent. The code-snippet for computing tangent stiffness using reverse-mode differentiation is given in Section 3.2. This method explicitly computes and stores the local tangent stiffness in a matrix form which then must be multiplied with the incremental strains $\delta\bm{e}$ to get the incremental stresses $\delta\bm{\sigma}$. This results in extra memory usage and two operations. {listing} {minted}[frame=lines, framesep=1mm, baselinestretch=0.7, fontsize=, ]python # create a function to compute stresses compute_stresses = jax.jacrev(strain_energy)
# computing the function to compute local tangent modulii K as given in Eq 4 compute_local_tangent = jax.jacfwd(compute_stresses)
# compute the tangent modulii tangent_modulii = compute_local_tangent(strain)
# compute the incremental stresses as given in Eq 4 dsigma = jax.numpy.einsum(’ijkl, kl -¿ ij’, tangent_modulii, dstrain) Automatic differentiation to compute incremental stresses using forward-mode AD
Automatic differentiation and especially the JAX library provides two other methods, i.e. the jvp and linearize functions, to ease the calculation of directional derivatives. In both approaches, the local tangent stiffness can be represented as a push-forward function (or a linear operator) that maps incremental strains $\delta\bm{e}$ lying in the tangent space of $\bm{e}$ to incremental stresses $\delta\bm{\sigma}$ lying in the tangent space of the stress domain. This facilitates the direct computation of incremental stress as a Jacobian-vector product [28, 29]. The jvp function (an example code-snippet is shown in Section 3.2) internally constructs the push-forward map and then applies it to the incremental strains to compute the incremental stresses. Representing $\mathbb{K}$ as a push-forward map rather than storing it explicitly in matrix form can significantly reduce the memory usage of FFT-based methods. However, creating this push-forward map with each function call can prove computationally expensive, particularly when utilized in the inner solver, like the conjugate gradient, of a Newton-Krylov solver. {listing} {minted}[frame=lines, framesep=1mm, baselinestretch=0.7, fontsize=, ]python # create a function to compute stresses compute_stresses = jax.jacrev(strain_energy)
# computing the incremental stresses as given in Eq 4 dsigma = jax.jvp(compute_stresses, (prev_strain,), (dstrain,))[1] Automatic differentiation to compute incremental stresses using jvp function
A computationally more efficient approach involves computing the local tangent stiffness around a fixed point, such as $\bm{e}$ from the previous converged step, within the Newton-Raphson scheme. This local stiffness can then be used to calculate incremental stresses during iterations in the inner solver. We utilize the JAX linearize function to generate the push-forward map for $\mathbb{K}$ at a specified strain point and store it as a function, which is then utilized to compute the incremental stresses. An example implementation is given in Section 3.2. {listing} {minted}[frame=lines, framesep=1mm, baselinestretch=0.7, fontsize=, ]python # create a function to compute stresses compute_stresses = jax.jacrev(strain_energy)
# compute the tangent stiffness function linearized around previous strain f_jvp = jax.linearize(compute_stresses, prev_strain)[1]
# compute the incremental stresses as given in Eq 4 dsigma = f_jvp(dstrain) Automatic differentiation to compute incremental stresses using linearize function
We examine the computational efficiency of these three methods to determine the incremental stresses from tangent stiffness on two scenarios: calculating incremental stresses for a single pixel (see Figure 2a) and for a grid of pixels (see Figure 2b). We consider a linear elastic material law for which the strain energy is given as $\psi=\lambda\mathrm{tr}(\bm{\varepsilon})^{2}+\mu\mathrm{tr}(\bm{\varepsilon}:\bm{\varepsilon})$ where $\lambda$ and $\mu$ are Lame’s parameters. For reference, we show the computational time when the local stiffness tangent and the incremental stresses are computed from the analytical expression. We observe that the calculation time for a pixel in forward or reverse mode, which are the two options for the first approach, is nearly as fast as using the analytical method (see Figure 2a). A small overhead occurs when using automatic differentiation due to the combined computation of derivatives and function values. However, employing the push-forward map via the linearize function or the jvp function increases the overall computation time for a single pixel (see Figure 2a). Although direct evidence for this cost increase is unavailable, we speculate that it results from constructing the push-forward map. When considering the computation time of the incremental stresses on a pixel grid (e.g. $63^{2}$, $127^{2}$, $255^{2}$), we observe that all the tested methods are on par with the analytic approach (see Figure 2b). Our results show that the jvp and linearize functions are slightly faster than the forward or reverse mode differentiation.
Figure 2: Computational efficiency of AD methods for the computation of incremental stresses from tangent stiffness. (a) Computational time for calculation on a single grid point. We consider a single function evaluation. (b) Computational time for a varying number of grid points. We compute the incremental stresses simultaneously overall the grid points.
3.3 Automatic computation of homogenized stiffness tensor